The sum of the present ages of Eric and his father is 65 years. In 8 years, his father will be twice as old as Eric will be at that time. Find their present ages.
whats erics age? and whats erics dads age?

Answer by scott8148(6628)

"The sum of the present ages of Eric and his father is 65 years" ___ E + f = 65 ___ f = 65 - E

"In 8 years, his father will be twice as old as Eric will be at that time" ___ f + 8 = 2(E + 8)

substituting ___ (65 - E) + 8 = 2E + 16 ___ 57 = 3E

solve for E, then substitute back to find f